Well, actully its pre-algebra, but anyways.
x/4 - 15/1 = -7
How do I do this? My teacher tries to explain, but it doesnt help at all, I don't understand.

Found 2 solutions by Alan3354, ReadingBoosters:
Answer by Alan3354(69443)

Well, actully its pre-algebra, but anyways.
x/4 - 15/1 = -7
How do I do this? My teacher tries to explain, but it doesnt help at all, I don't understand.
-------------
Do what?
You didn't say what you want to do.

Answer by ReadingBoosters(3246)

Starting with
x/4 - 15/1 = -7
15/1 is the same as 15
x/4 - 15 = -7
Add 15 to both sides
x/4 - 15 + 15 = -7 + 15
x/4 = 8
Multiply both sides by 4
(x/4)(4) = 8(4)
x = 32
Alternatively, another method, especially if the fraction had a denominator other than 1:
x/4 - 15/1 = -7
Multiply both sides by 4
(x/4)(4) - (15/1)(4) = -7(4)
x - 60 = -28
Add 60 to both sides
x - 60 + 60 = -28 + 60
x = 32
.......................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.